### v1

# When Is Diversity Rewarded in Cooperative Multi-Agent Learning?

Abstract
The success of teams in robotics, nature, and society often depends on the division of labor among diverse specialists; however, a principled explanation for when such diversity surpasses a homogeneous team is still missing. Focusing on multi-agent task allocation problems, our goal is to study this question from the perspective of reward design: what kinds of objectives are best suited for heterogeneous teams? We first consider an instantaneous, non-spatial setting where the global reward is built by two generalized aggregation operators: an inner operator that maps the $N$ agents’ effort allocations on individual tasks to a task score, and an outer operator that merges the $M$ task scores into the global team reward. We prove that the curvature of these operators determines whether heterogeneity can increase reward, and that for broad reward families this collapses to a simple convexity test. Next, we ask what incentivizes heterogeneity to emerge when embodied, time-extended agents must learn an effort allocation policy. To study heterogeneity in such settings, we use multi-agent reinforcement learning (MARL) as our computational paradigm, and introduce Heterogeneous Environment Design (HED), a gradient-based algorithm that optimizes the parameter space of underspecified MARL environments to find scenarios where heterogeneity is advantageous. Experiments in matrix games and an embodied Multi-Goal-Capture environment show that, despite the difference in settings, HED rediscovers the reward regimes predicted by our theory to maximize the advantage of heterogeneity, both validating HED and connecting our theoretical insights to reward design in MARL. Together, these results help us understand when behavioral diversity delivers a measurable benefit.
Figure 1: We study and categorize what reward structures lead to the need for behavioral heterogeneity in multi-agent multi-task environments.
1 Introduction
Collective systems—from robot fleets to insect colonies—tend to adopt one of two fundamental structures: a uniform shared blueprint or a set of distinct, specialized roles. In multi-agent learning, this dichotomy is reflected in the design choice between behavioral homogeneity (where all agents share one policy) and heterogeneity (where each agent learns its own) bettini2023hetgppo ; bettini2023snd ; rudolph2021desperate . Although diversity unlocks role specialization and asymmetric information use, it also introduces extra coordination cost, representation overhead, and learning complexity chenghao2021celebrating . This trade-off leads us to ask: under what conditions will heterogeneous agents actually outperform the best homogeneous baseline?
A natural setting to study this question in is multi-agent task allocation, where $N$ agents allocate effort across $M$ concurrent tasks. A homogeneous team is forced to adopt a single allocation vector (e.g., every agent spends $0.75$ of its effort on task A and $0.25$ on task B), whereas a heterogeneous team allows agents to specialize differently. This allocation setting underlies multi-agent problems such as coverage and load-balancing Gupta2017 ; Lowe2017 . In this work, we relate it to heterogeneity in cooperative navigation (Sec. 5), Colonel Blotto games, and level-based foraging environments (App. D) Roberson2006 ; noel2022reinforcementcolonelblotto ; papoudakis2021benchmarking_multilevelforaging ; terry2021pettingzoo .
[…]
To study heterogeneity in MARL settings not covered by our theoretical analysis, we develop Heterogeneous Environment Design (HED), a gradient-based co-design algorithm that learns environment parameters $\theta$ in under-specified tasks via backpropagation, leveraging differentiable simulation to find parameterizations that either maximize or minimize $\Delta R$. In our experiments $\theta$ parameterizes the reward structure, i.e., the inner- and outer-aggregators, but any differentiable feature of the environment (e.g., obstacle placements) can be optimized. Maximizing the heterogeneity gain through HED allows us to discover previously unknown types of tasks where heterogeneity is essential; minimizing the gain steers an underspecified task toward regimes where homogeneous policies are sufficient, which may be desirable, e.g., when parameter sharing is preferred for learning efficiency (Sec. 4).
We evaluate the validity of our theoretical insights, and HED, in simulation. To this end, we design single-shot and long-horizon reinforcement learning environments whose reward structure instantiates the kinds of aggregation operators we studied. First, in a continuous and a discrete matrix game, we test reward structures based on all nine possible combinations of $\{\min,\mathrm{mean},\max\}$, and find that the heterogeneity gains that result from the agents’ learned policies match our theoretical predictions: concave outer aggregators and convex inner aggregators benefit heterogeneous teams. Next, we test the same aggregators in an embodied, partially observable Multi-Goal-Capture environment, and find that our theory also transfers to this long-horizon MARL setting. Importantly, we also show that reward structures that maximize heterogeneity are meaningful and practically useful. Finally, we find that the empirical heterogeneity gain disappears as the richness of agents’ observations is increased, recovering the finding that rich observations allow agents with identical policy networks to be behaviorally heterogeneous bettini2023hetgppo ; leibo2019malthusian .
We then turn to HED. Across two parameterizable families of aggregators (Softmax and Power-Sum), we show that, despite running on spatial and long-horizon environments, HED rediscovers the reward regimes predicted by our curvature theory to maximize the heterogeneity gain, both validating HED and confirming the connection between our theoretical insights and MARL reward design (Sec. 5).
1.1 Related Works
Behavioral heterogeneity—where capability-identical agents learn distinct policies—can markedly improve exploration, robustness, and reward bettini2023hetgppo . Yet heterogeneity reduces parameter sharing and thus sample-efficiency, so a core practical question is when its benefits outweigh that cost. Existing MARL methods typically adopt one of two poles: endowing each agent with its own network, or enforcing parameter sharing so that all agents follow a single policy gupta2017cooperative ; rashid2018qmix ; foerster2018counterfactual ; kortvelesy2022qgnn ; sukhbaatar2016learning . A substantial body of work explores the efficiency–diversity trade-off christianos2021scaling ; fu2022revisiting by interpolating between these extremes—e.g. injecting agent IDs into the observation foerster2016learning ; gupta2017cooperative , masking different subsets of shared weights NEURIPS2024_274d0146 , sharing only selected layers chenghao2021celebrating , pruning a shared network into agent-specific sub-graphs kim2023parameter , or producing per-agent parameters with a hypernetwork tessera2024hypermarl . Further, several methods for promoting behavioral diversity in MARL have been proposed, such as: conditioning agents’ policies on a latent representation wang2020roma , decomposing and clustering action spaces wang2021rode , dynamically grouping agents to share parameters yang2022ldsa , applying structural constraints to the agents’ policies bettini2024controlling , or by additional tasks and intrinsic rewards that maximize diversity chenghao2021celebrating ; jaques2019social ; wang2019influence ; jiang2021emergence ; mahajan2019maven ; liu2023contrastive ; liu2024interaction ; li2024learning . While these studies demonstrate how to obtain diversity, they all presume tasks where heterogeneity is already advantageous; what is still missing—and what we supply here—is a principled characterization of which reward structures create that incentive in the first place.
Classic resource–allocation settings, in which a team must divide finite effort among simultaneous objectives, are a central proving ground for cooperative MARL. In robotics, potential-field and market-based learning are the dominant tools for coverage, exploration, and load-balancing tasks Gupta2017 ; Lowe2017 . Game-theoretic analysis and, recently, MARL, play the same role in discrete counterparts such as Colonel-Blotto contests, where players decide how to spread forces over several “battlefields” Roberson2006 ; noel2022reinforcementcolonelblotto . Embodied benchmarks like level-based foraging are heavily studied in MARL, and expose the tension between uniform and specialized effort allocations papoudakis2021benchmarking_multilevelforaging . The survey of BasarOverview highlights how cooperative performance is governed by the shape of the shared reward and the equilibria it induces. Our contribution sharpens this perspective: we prove that the curvature of nested aggregation operators characterizes when heterogeneous allocations dominate homogeneous ones, and introduce algorithmic tools for further exploring settings where diversity is needed.
Co-design is a paradigm where agent policies and their mission or environment are simultaneously optimized gao2024codesign . Our HED algorithm is related to PAIRED dennis2020emergent , a method which automatically designs environments in a curriculum such that an antagonist agent succeeds while the protagonist agent fails. This makes it so that resulting environments are challenging enough without being unsolvable. Similarly, HED designs environments that are advantageous to heterogeneous teams, while disadvantaging homogeneous teams. The key differences are: (1) the environment designer uses direct regret gradient backpropagation via a differentiable simulator instead of RL; this enables higher efficiency by directly leveraging all the environment gradient data available during collection while preventing RL-related issues identified in subsequent works jiang2021replayguided ; parker-holder2021that such as exploration inefficiency and the need for a reward signal; and (2), the protagonist and antagonist are independent multi-agent teams instead of single agents.
2 Formal Setting: Multi-Agent Task Allocation
Consider a set of $N$ agents and $M$ tasks. Each agent $i\in\{1,\ldots,N\}$ allocates effort among the tasks according to the budget constraints: $r_{i1},r_{i2},\ldots,r_{iM}\geq 0\text{ with }\sum_{j=1}^{M}r_{ij}\leq 1$. We can consider both continuous allocations ($r_{ij}$ can be any real number) and discrete allocations ($r_{ij}$ restricted to some finite set of options), with most results in this work focusing on the continuous case. We collect all agents’ allocations into an $N\times M$ matrix: $\mathbf{A}=[r_{ij}]$ [^1] All results in this work can be extended to the case where $r_{i1},r_{i2},\ldots,r_{iM}\geq B_{min}\text{ and }\sum_{j=1}^{M}r_{ij}\leq B_{max}$ for some arbitrary $B_{min},B_{max}\in\mathbb{R}$..
[…]
A homogeneous strategy is one where all agents have the same allocation (i.e., devote the same amount of effort to a given task $j$): $r_{ij}=c_{j}\forall\,i,j$. In this case, the allocation matrix $\mathbf{A}$ consists of identical rows. We define $R_{\mathrm{hom}}=\max_{(c_{1},\ldots,c_{M})\in\Delta_{\!\leq}^{M-1}}\;R\Bigl{(}\mathbf{A}\Bigr{)}$ where $\Delta_{\!\leq}^{M-1}=\{(c_{1},\ldots,c_{M})\mid c_{j}\geq 0,\;\sum_{j}c_{j}\leq 1\}$ is the closed unit simplex. A heterogeneous strategy allows each agent $i$ to choose any $(r_{i1},\ldots,r_{iM})\in\Delta_{\!\leq}^{M-1}$ independently. Then $R_{\mathrm{het}}=\max_{\mathbf{A}\in(\Delta_{\!\leq}^{M-1})^{N}}\;R\Bigl{(}\mathbf{A}\Bigr{)}.$ We define the heterogeneity gain as: $\Delta R=R_{\mathrm{het}}-R_{\mathrm{hom}}.$ This quantity measures how much greater the overall reward can be when agents are allowed to specialize differently across tasks, compared to when they must behave identically. Characterizing when $\Delta R>0$ is our main focus in this work.
In MARL, the value $r_{ij}$ represents the ‘local rewards’ agents receive based on task performance, and the aggregate reward $R(\mathbf{A})$ can represent: (i) the payoff of a one-shot effort-allocation game, (ii) the return or sparse terminal reward of an episode, or (iii) the stepwise reward, giving the discounted return $\sum_{t=0}^{T}\gamma^{t}\,R\bigl{(}\mathbf{A}_{t}\bigr{)}$ for a sequence $\bigl{(}\mathbf{A}_{t}\bigr{)}_{t=1,\ldots T}$ of allocations [^2] We can generalize (iii) by considering a sequence $(R_{t})_{t=1\ldots,T}$ of allocation returns that varies per time step.. A positive gain $\Delta R>0$ implies heterogeneous agents can outperform homogeneous ones if the heterogeneous agents can always attain the optimal allocation; if not (as in our Multi-Goal-Capture experiment), this is evidence of an advantage to heterogeneity, but not a formal guarantee.
App. H contains concrete examples of generalized aggregators. In Sec. 5, we study two MARL environments whose reward structure conforms to our setting: a one-shot task allocation game and an embodied Multi-Goal-Capture environment, which can be seen as an extension of various cooperative navigation environments (see terry2021pettingzoo ). Finally, in App. D, we show how the heterogeneity gain of two well-known environments from the literature–Team Colonel Blotto games noel2022reinforcementcolonelblotto and level-based foraging papoudakis2021benchmarking_multilevelforaging –can be studied using the theory we develop in this work.
3 Analysis
Focusing on continuous allocations, we ask what properties of aggregators guarantee $\Delta R>0$. We draw on the concept of Schur-convexity. Schur-convex functions can be understood as generalizing symmetric, convex aggregators: every convex and symmetric function is Schur-convex, but a Schur-convex function is not necessarily convex roberts1974convex ; peajcariaac1992convex . Proofs for all results are available in App. F.
Since both the outer aggregator $U$ and the task-level aggregators $T_{j}$ are non-decreasing, an optimal effort allocation will always have each agents’ efforts summing to $1$. Hence, from here on, we assume without loss of generality that $\sum_{j=1}^{M}r_{ij}=1$. We call such allocations admissible.
Definition 3.1 (Majorization).
Let $x=(x_{1},\dots,x_{N})$ and $y=(y_{1},\dots,y_{N})$ be two vectors in $\mathbb{R}^{N}$ such that $\sum_{i=1}^{N}x_{(i)}=\sum_{i=1}^{N}y_{(i)}$. Let $x_{(1)}\geq x_{(2)}\geq\cdots\geq x_{(N)}$ and $y_{(1)}\geq y_{(2)}\geq\cdots\geq y_{(N)}$ be the components of $x$ and $y$ sorted in descending order. We say that $x$ majorizes $y$ (written $x\succ y$) if $\sum_{i=1}^{k}x_{(i)}\geq\sum_{i=1}^{k}y_{(i)}$ for $k=1,2,\dots,N-1,N.$
Definition 3.2 (Schur-Convex Function).
A symmetric function $f:\mathbb{R}^{N}\to\mathbb{R}$ is Schur-convex if for any two vectors $x,y\in\mathbb{R}^{N}$ with $x\succ y$, we have $f(x)\geq f(y).$ If the inequality is strict whenever $x$ and $y$ are not permutations of each other, then $f$ is said to be strictly Schur-convex. Similarly, $f$ is Schur-concave if $f(x)\leq f(y)$ whenever $x\succ y$.
Intuitively, $x\succ y$ means one can obtain $y$ from $x$ by repeatedly moving mass from larger to smaller coordinates, thereby making the vector more uniform. Schur-convexity is then a statement on a function’s curvature: $f$ is Schur-convex if it increases with inequality, or is Schur-concave if it increases with uniformity. We show here a connection between Schur-convexity (concavity) and $\Delta R$.
Call an allocation matrix $\mathbf{A}$ trivial if there exists a task $j^{\star}$ such that every agent allocates its entire budget to that task, i.e. $r_{ij^{\star}}=B_{\max}$ and $r_{ij}=0\ \forall i,\;\forall j\neq j^{\star}$; otherwise $A$ is non-trivial. Then:
Theorem 3.1 (Positive Heterogeneity Gain via Schur-convex Inner Aggregators).
Let $N,M\geq 2$, and assume that (i) each task-level aggregator $T_{j}$ is strictly Schur-convex and (ii) the outer aggregator $U$ is coordinate-wise strictly increasing. Then either all admissible optimal homogeneous allocations are trivial, or $\Delta R>0$.
If the task-level aggregator is instead Schur-concave, we can show there is no heterogeneity gain:
Theorem 3.2 (No Heterogeneity Gain via Schur-concave Inner Aggregators).
[…]
We see that Schur-convexity of the inner aggregator produces $\Delta R>0$, whereas Schur-concavity implies $\Delta R=0$. Analyzing the outer aggregator $U$ is trickier, because it acts on task-score vectors $\bigl{(}T_{1}(\mathbf{a}_{1}),\dots,T_{M}(\mathbf{a}_{M})\bigr{)}$ whose sum $\sum_{i=1}^{M}T_{i}(\mathbf{a}_{i})$ may vary, so majorization is not directly applicable. However, we can extend our analysis to $U$ if our inner aggregators are normalized to keep the sum constant: $\sum_{i=1}^{M}T_{i}(\mathbf{a}_{i})=C$ for any admissible allocation. Assuming this, we can invoke majorization again, and the relationship between convexity and $\Delta R$ reverses: if the outer aggregator $U$ is Schur-convex, the heterogeneity gain vanishes. Let us prove this.
Theorem 3.3 (No Heterogeneity Gain for Schur-Convex $U$ with Constant-Sum Task Scores).
Let $N,M\geq 2$. Suppose that for any admissible allocation $\mathbf{A}$, (i) every task score is non-negative, and obeys $T_{i}(0,\ldots,0)=0$, and (ii) the sum of task score is always $\sum_{j=1}^{M}T_{j}\bigl{(}\mathbf{a}_{j}\bigr{)}\;=\;C$. If $U$ is strictly Schur-convex function, then $\Delta R=0.$
Sum-Form Aggregators. In App. E, we show that the above results reduce to a simple convexity test for sum-form aggregators: a broad class of aggregators that describes most reward structures we consider in this work. This makes testing whether $\Delta R>0$ a simple computation in many cases.
A core topic of this work is reward design: how can we craft team objectives that either advantage or disadvantage behavioral diversity? To do this, it is helpful to first identify an appropriate search space. Our theoretical analysis enables us to narrow down this search space, and focus on aggregators whose curvature can be parametrized. Many family of aggregator functions $\{\,f_{t}(\cdot)\}_{t\in\mathbb{R}}$ can be parametrized by a scalar $t$ which controls whether the aggregation is Schur-convex or Schur-concave, and how strongly it penalizes (or favors) inequalities among the components. For example, the softmax aggregator $\sum_{i=1}^{N}\frac{\exp\bigl{(}t\cdot r_{ij}\bigr{)}}{\sum_{\ell=1}^{N}\exp\bigl{(}t\cdot r_{\ell j}\bigr{)}}$ is parametrized by its temperature, $t$, transitioning from being strictly Schur-concave when $t<0$ to strictly Schur-convex when $t>0$. We can define a space of reward functions by selecting both the task scores and outer aggregator to be softmax functions: let $T_{j}(\mathbf{A})=\sum_{i=1}^{N}\frac{\exp\bigl{(}t\cdot r_{ij}\bigr{)}}{\sum_{\ell=1}^{N}\exp\bigl{(}t\cdot r_{\ell j}\bigr{)}}\;r_{ij},$ and let $U\bigl{(}T_{1}(\mathbf{a}_{1}),\ldots T_{M}(\mathbf{a}_{m}))=\sum_{j=1}^{M}\frac{\exp\bigl{(}\tau\cdot T_{j}(\mathbf{A})\bigr{)}}{\sum_{\ell=1}^{M}\exp\bigl{(}\tau\cdot T_{\ell}(\mathbf{A})\bigr{)}}\;T_{j}(\mathbf{A})$, where $t,\tau\in\mathbb{R}$ parametrize the inner and outer aggregators, respectively. The heterogeneity gain is then dependent on $t$ and $\tau$. As a case study, we derive lower bounds on $\Delta R$ when $N=M$ in Thm. 3.4. We conjecture the lower bound of Thm. 3.4 is actually the exact value of $\Delta R(t,\tau;N)$, but we were unable to prove this. Fig. 2 (righthand side) plots the heterogeneity gains when $N=M=2$.
Theorem 3.4 (Exact softmax heterogeneity gain for $N=M$).
Assume $N=M\geq 2$, and let $\sigma(t,N):=\frac{e^{t}}{e^{t}+N-1}$. The heterogeneity gain for softmax aggregators (i) equals $\Delta R(t,\tau;N)=0$ when $t\leq 0$; (ii) is bounded below by $\sigma(t,N)-\dfrac{1}{N}$ when $t>0,\tau\leq 0$; and (iii) is bounded below by $\max\!\bigl{\{}\sigma(t,N)-\sigma(\tau,N),0\bigr{\}}$ when $t>0,\tau\geq 0.$
Tab. 3 contains more examples of aggregation operators parameterized by $t$. These families provide a search space for potential reward functions, allowing us to sweep smoothly from $\Delta R=0$ to $\Delta R>0$ reward regimes. As $t\to\pm\infty$, most such aggregators converge to either $\min$ or $\max$, and often reduce to the arithmetic mean for certain parameter choices, motivating us to ask what the heterogeneity gain is when the outer and inner aggregator belong to the set $\{\min,\text{mean},\max\}$. These aggregators are of special interest, since “$\min$” can be seen as a “maximally” Schur-concave function, “$\max$” can be seen as a “maximally” Schur-convex function, and “mean” is both Schur-convex and Schur-concave. Hence, it is worth asking what the heterogeneity gain is when the outer and inner aggregator belong to the set $\{\min,\text{mean},\max\}$. We derive these gains in two cases: continuous allocations where $r_{ij}\in[0,1]$, and discrete effort allocations where $r_{ij}\in\{0,1\}$. The results are summarized in Fig. 2, lefthand side (formal derivation available in App. F).
Discrete and continuous heterogeneity gains
$T=\min$ $T=\text{mean}$ $T=\max$ Outer $U=\min$ $\Delta R_{\mathrm{F}}$ 0 0 $(M-1)/M$ $\Delta R_{\mathrm{D}}$ 0 $\lfloor N/M\rfloor/N$ $\mathbf{1}_{\{N\geq M\}}$ Outer $U=\text{mean}$ $\Delta R_{\mathrm{F}}$ 0 0 $(M-1)/M$ $\Delta R_{\mathrm{D}}$ 0 0 $(\min\{M,N\}-1)/M$ Outer $U=\max$ $\Delta R_{\mathrm{F}}$ 0 0 0 $\Delta R_{\mathrm{D}}$ 0 0 0
Figure 2: Left: Discrete ($\Delta R_{\mathrm{D}}$) and continuous-allocation ($\Delta R_{\mathrm{F}}$) heterogeneity gains for all $U,T\!\in\!\{\min,\text{mean},\max\}$. The indicator $\mathbf{1}_{\{N\geq M\}}$ equals 1 if $N\geq M$ and 0 otherwise. Right: We plot the parametrized heterogeneity gains $\Delta R(t,\tau;N)$ when $U$ and $T$ are soft-max aggregators.
4 Heterogeneous Environment Design (HED)
In complex scenarios where theory might be less applicable, we study heterogeneity through algorithmic search. We consider the setting of a Parametrized Dec-POMDP (PDec-POMDP, defined in App. K). A PDec-POMDP represents a Dec-POMDP oliehoek2016concise , where the observations, transitions, or reward are conditioned on parameters $\theta$. Hence, the return obtained by the agents, $G^{\theta}(\mathbf{\pi})$, can be differentiated with respect to $\theta$: $\nabla_{\theta}G^{\theta}(\mathbf{\pi})=\frac{\partial}{\partial_{\theta}}G^{\theta}(\mathbf{\pi})$. In particular, computing this gradient in a differentiable simulator allows us to back-propagate through time and optimize $\theta$ via gradient ascent [^3] Although the same approach can train policies xu2022accelerated ; song2024learning , HED instead optimizes environment parameters and policies separately, using standard zeroth-order policy-gradient methods, to avoid being trapped in local minima..
Algorithm 1 Heterogeneous Environment Design (HED).
0:  Environment parameters $\theta$, environment learning rate $\alpha$, heterogeneous agent policy $\mathbf{\pi}_{\mathrm{het}}$, homogeneous agent policy $\mathbf{\pi}_{\mathrm{hom}}$
1:  for $i$ in iterations do
[…]
5:     if train_env($i$then
6:        $\theta\leftarrow\theta+\alpha\nabla_{\theta}\mathrm{HetGain}^{\theta}$ train environment via backpropagation}
7:     if train_agents($i$then
[…]
9:  final environment configuration $\theta$, policies $\mathbf{\pi}_{\mathrm{het}},\mathbf{\pi}_{\mathrm{hom}}$
We now consider the problem of learning the environment parameters $\theta$ to maximize the empirical heterogeneity gain. The empirical heterogeneity gain is defined as the difference in performance between heterogeneous and homogeneous teams in a given PDec-POMDP parametrization. Heterogeneous agents each learn different parameters, whereas homogeneous agents learn the same policy. We denote their policies as $\mathbf{\pi}_{\mathrm{het}}$ and $\mathbf{\pi}_{\mathrm{hom}}$. Then, we can simply write the gain as: $\mathrm{HetGain}^{\theta}(\mathbf{\pi}_{\mathrm{het}},\mathbf{\pi}_{\mathrm{hom}})=G^{\theta}(\mathbf{\pi}_{\mathrm{het}})-G^{\theta}(\mathbf{\pi}_{\mathrm{hom}})$, representing the return of heterogeneous agents minus that of homogeneous agents on environment parametrization $\theta$. HED, shown in Alg. 1, learns $\theta$ by performing gradient ascent to maximize the gain: $\theta\leftarrow\theta+\alpha\nabla_{\theta}\mathrm{HetGain}^{\theta}(\mathbf{\pi}_{\mathrm{het}},\mathbf{\pi}_{\mathrm{hom}})$. The environment and the agents are trained in an iterative process. At every training iteration, HED collects roll-out batches in the current environment $\theta$ for both heterogeneous and homogeneous teams, computing the heterogeneity gain on the collected data. Then, it updates $\theta$ to maximize the heterogeneity gain. Finally, to train the agents, it uses MARL, with any on-policy algorithm (e.g., MAPPO yu2022surprising ). The functions train_env and train_agents determine when to train each of the components in HED. In particular, we consider two possible training regimes: (1) alternated: where HED performs cycles of $x$ agent training iterations followed by $y$ environment training iterations and (2) concurrent: where agents train at every iteration and the environment is updated every $x$ iterations. Note that by performing descent instead of ascent, HED can also be used to minimize the heterogeneity gain, which may be useful in homogeneous systems.
5 Experiments
To empirically ground our theoretical analysis, we conduct a three-stage experimental study in cooperative MARL. We first analyze a one-step, observation-free matrix game in which each agent allocates effort $r_{ij}$ over $M$ tasks, and consider reward structures defined by aggregator pairs $U,T\in\{\min,\mathrm{mean},\max\}$. We find that the agents’ learned policies recover the exact heterogeneity gains derived in the theory (Fig. 2). Next, we transfer the same reward structures into an embodied, time-extended Multi-Goal-Capture environment. We show that our curvature theory continues to be predictive in this setting as well, and discuss interesting learning dynamics emerging from the agents’ embodiment. We also expose a trade-off between the heterogeneity gain and agents’ observations, showing that when homogeneous agents can access richer observations, they can learn incrementally heterogeneous behaviors by leveraging behavioral typing and thus minimizing the gain: this highlights the difference between neural and behavioral heterogeneity bettini2023hetgppo .
Finally, to study HED, we parametrize the reward structure of Multi-Goal-Capture using either parametrized Softmax or Power-Sum aggregators (App. H), and run HED to learn parameterizations that maximize the heterogeneity gain. HED learns the theoretically optimal aggregator instantiations: one the one hand, this validates HED’s effectiveness at discovering heterogeneous missions, and on the other hand, showcases the generalizability of our theoretical insights to embodied environments. Implementation details and visualizations are available in App. C and App. J, respectively.
We consider a one-step observationless matrix game where $N$ agents need to choose between $M$ tasks. Their actions are effort allocations $r_{ij}$ with $r_{ij}\geq 0$,$\sum_{j}r_{ij}=1$, composing matrix $\mathbf{A}$. With aggregators taken from the set $U,T\in\{\min,\mathrm{mean},\max\}$, our goal is to empirically confirm the heterogeneity gains derived in the theory in a learning context. Each time the game is played, all agents obtain the global reward $R(\mathbf{A})$ computed through the double aggregator. We consider two setups: (1) Continuous ($r_{ij}\in\mathbb{R}_{0\leqslant x\leqslant 1}$): agents can distribute their efforts across tasks, (2) Discrete ($r_{ij}\in\{0,1\}$): agents choose only one task. We train with $N=M=4$ for 12 million steps. The evolution of the heterogeneity gains over training is shown in Fig. 6. The final results match exactly the theoretical predictions in the Fig. 2 table and our curvature theory: concave outer and convex inner aggregators favor heterogeneity. Additional results and details are in App. I.
Figure 3: Heterogeneity gain for the discrete and continuous matrix games with $N=M=4$ over training iterations. We report mean and standard deviation after 12M frames over 9 random seeds. The final results match the theoretical predictions in the Table of Fig. 2.
Next, we investigate a time-extended, embodied scenario called Multi-Goal-Capture, based on multi-goal navigation missions terry2021pettingzoo . Our goal is to show that the results obtained in the matrix game still hold in this more complex setting and that aggregators that maximize the heterogeneity gain encode a practically useful global objective.
In Multi-Goal-Capture, agents need to navigate to goals (visualization availabile in App. J). Each agent observes the relative position to the goals, and agent actions are continuous 2D forces that determine their direction of motion. The entries $r^{t}_{ij}$ of matrix $\mathbf{A}^{t}$ at time $t$ represent the local reward of agent $i$ towards goal $j$, computed as $r^{t}_{ij}=\left(1-d^{t}_{ij}/\sum_{j=1}^{M}d^{t}_{ij}\right)/(M-1),$ where $d^{t}_{ij}$ is the distance between agent $i$ and goal $j$. This makes it so that $\sum_{j=1}^{M}r^{t}_{ij}=1$ and $r^{t}_{ij}\geqslant 0$. At each step, the agents receive the global reward $R(\mathbf{A}^{t})$, again with aggregators $U,T\in\{\min,\mathrm{mean},\max\}$. Different aggregator choices will yield different global objectives. For example, $U=\max,T=\max$ implies “at least one agent should go to at least one goal”; $U=\max,T=\min$ implies “all agents should go to the same goal”, and so on. $U=\min,T=\max$, a concave-convex setting shown by our theory to favor heterogeneity, implies “each agent should go to a different goal and all goals should be covered” which is a natural goal for this scenario. This is because $T=\max$ encodes a task that needs just one agent to be completed (e.g., find an object), while $U=\min$ encodes that all tasks should be attended (i.e., agents need to diversify their choices).
(a) Gain for different aggregators.
(b) Gain w.r.t. observability when $U=\min,T=\max$.
Figure 4: Heterogeneity gain for Multi-Goal-Capture throughout training. We report mean and standard deviation for 30 million training frames over 9 (left), 4 (right) random seeds.
After 30M environment frames (Fig. 4(a)) the empirical heterogeneity gains differ, numerically, from those of the static matrix-game because agents now realize their allocations $r_{ij}$ through time-extended motion. Nonetheless, our curvature theory still predicts when there is a heterogeneity gain (Fig. 2): it is positive only for the concave–convex pairs $U=\min,\,T=\max$ and $U=\mathrm{mean},\,T=\max$. The heterogeneity gain is smaller in the latter case because learning dynamics matter: with $U=\min,\,T=\max$ the best homogeneous policy is unique–every agent must steer to the midpoint between the two goals–so homogeneous learners seldom find it, leaving room for heterogeneous policies to excel (see App. J). By contrast, $U=\mathrm{mean},\,T=\max$ admits a continuum of good homogeneous policies, which homogeneous teams execute more easily. For $U=\max,\,T=\min$ and $U=\max,\,T=\mathrm{mean}$ the empirical heterogeneity gap is actually negative: this is because the reward peaks only when all agents choose the same goal, a coordination that heterogeneous teams learn more slowly, so they lag within the fixed training budget. Crucially, while additional training will eventually result in $\Delta R=0$ in the latter case, the positive heterogeneity gains we report are theoretically irreducible, pinpointing scenarios where behavioral diversity is needed.
Observability-Heterogeneity Trade-Off: We now crystallize the relationship between environment observability and empirical heterogeneity gains. It is well known that neurally homogeneous agents (i.e., sharing the same neural network) can emulate diverse behavior by conditioning on the input context (behavioral typing) leibo2019malthusian . This can be achieved by naively appending the agent index to its observation gupta2017cooperative or by providing relevant observations that allows the agents to infer their role bettini2023hetgppo . Behavioral typing is impossible in matrix games, as these games are observationless. However, it is possible in more complex games, such as our Multi-Goal-Capture scenario. We augment agents in the positive gain scenario ($U=\min,T=\max$) with a range sensor, providing proximity readings for other agents within a radius. In 4(b), we show that the heterogeneity gain decreases as the agent visibility increases (higher sensing radius). This is because, with a higher range, homogeneous agents can sense each other and coordinate to pursue different goals. This result highlights the tight interdependence between the heterogeneity gain and agents’ observations.
We apply HED to Multi-Goal-Capture, and ask whether it finds the same aggregator parameterizations predicted by our theory to maximize the heterogeneity gain. We turn the environment into a PDec-POMDP by parameterizing the reward as ${R}^{\theta}(\mathbf{A}^{t})=\bigoplus_{j=1}^{M}\!{}^{\theta}\bigoplus_{i=1}^{N}\!{}^{\theta}r^{t}_{ij}$, with parametrized inner and outer aggregators $U^{\theta}=\bigoplus^{\theta},T^{\theta}=\bigoplus^{\theta}$. Our goal is to learn the parameters $\theta=(\tau_{1},\tau_{2})$, parametrizing $T$ and $U$ respectively, that maximize the heterogeneity gain. We consider two parametrized aggregators from Tab. 3: Softmax and Power-Sum. Softmax: We parameterize both $U^{\theta}$ and $T^{\theta}$ using Softmax. We initialize $\tau_{1}=\tau_{2}=0$, so $U$ and $T$ are initially mean, and run HED. In 5(a), we show that, to maximize the heterogeneity gain, HED learns to maximize $\tau_{1}$, making the inner aggregator $T$ Schur-convex, while minimizing $\tau_{2}$, making the outer aggregator $U$ Schur-concave. The result is exactly the reward function predicted by our theory to maximize the gain. Power-Sum: We parametrize both aggregators with Power-Sum. We initialize both functions to $\tau_{1}=\tau_{2}=1$, representing sum, and run HED. We constrain $\tau_{1,2}\in[0.3,6]$ to stabilize learning. In 5(b), we show that HED learns to maximize $\tau_{1}$, making $T$ Schur-convex, while minimizing $\tau_{2}$, making $U$ Schur-concave; again rediscovering the optimal parametrization our theory predicts. Hence, these results simultaneously validate HED and our curvature theory.
(a) HED with Softmax aggregators ($\tau\in\mathbb{R}$). $\tau\geq 0$ is Schur-convex; $\tau\leq 0$ is Schur-concave.
(b) HED with Power-Sum aggregators ($\tau\in[0.3,6]$). $\tau\geq 1$ is Schur-convex; $\tau\leq 1$ is Schur-concave.
Figure 5: HED results in Multi-Goal-Capture. The two leftmost columns report the evolution of aggregator parameters through training, while the rightmost column shows the obtained heterogeneity gain. This result empirically demonstrates that HED rediscovers the reward structure predicted by our theory to maximize the gain, making the inner aggregator convex, and the outer aggregator concave. We report mean and standard deviation for 90M training frames over 13 random seeds.
6 Discussion
This work introduces tools for both diagnosing and designing reward functions that incentivize heterogeneity in cooperative MARL. In task allocation settings, our theory shows that the advantage of behavioral diversity is a predictable consequence of reward curvature: if the inner aggregator is Schur-convex, amplifying inequality, and the outer aggregator is Schur-concave, amplifying uniformity, heterogeneous policies are strictly superior; reversing the curvature removes the benefit. Complementing this analysis, and covering settings where our theory doesn’t apply, the proposed HED algorithm automatically steers underspecified environments to either side of the diversity boundary, letting us encourage or suppress heterogeneity and providing a sandbox for studying its advantages. Together, these results help turn the choice of heterogeneity from an ad-hoc heuristic into a controllable design dimension, and help reconcile past mixed results on parameter sharing.
A key remaining open question concerns how the environment’s transition dynamics interact with reward curvature to shape heterogeneity gains. It is also important to consider how one can extend our findings to inseparable tasks. We expand on these open questions, and other limitations, in App. L.
Acknowledgments and Disclosure of Funding
This work is supported by European Research Council (ERC) Project 949940 (gAIa) and ARL DCIST CRA W911NF-17-2-0181. We gratefully acknowledge their support.
Appendix A Computational Resources Used
For the realization of this work, we have employed computational resources that have gone towards: experiment design, prototyping, and running final experiment results. Simulation and training are both run on GPUs, no CPU compute has been used. Results have been stored on the WANDB cloud service. We estimate:
  • 300 compute hours on an NVIDIA GeForce RTX 2080 Ti GPU.
  • 500 compute hours on an NVIDIA L40S GPU.
Simply verifying our results using the available code will take considerably less compute hours (around a day).
Appendix B Code and Data Availability
We attach the code in the supplementary materials. The code contains instructions on how to reproduce the experiments in the paper and dedicated YAML files containing the hyperparameters for each experiment presented. The YAML files are structured according to the HYDRA [45] framework which allows smooth reproduction as well as systematic and standardized configuration. We further attach all scripts to reproduce the plots in the paper from the experiment results.
Appendix C Implementation Details
For all experiments, we use the MAPPO MARL algorithm [44]. Environments are implemented in the multi-agent environment simulator VMAS [46], and trained using TorchRL [47]. Both the actor and critic are two-layer MLPs with 256 neurons per layer and Tanh activation. Further details, such as hyperparameter choices, are available in the attached code and YAML configuration files.
Appendix D Colonel Blotto & Level-Based Foraging
We describe how two well-known settings from the literature fit into our theoretical framework, and check what our theoretical results say about their heterogeneity gain.
D.1 Team Colonel Blotto (fixed adversary)
The Colonel Blotto game is a well-known allocation game studied in both game theory and MARL [7, 8]. It is used to model election strategies and other resource-based competitions. In the team variant with fixed adversary, $N$ friendly colonels (agents) each distribute a (fixed and equal) budget of troops $r_{ij}\!\geq 0,\sum_{j=1}^{M}r_{ij}=1$ across $M$ battlefields $j\in\{1,\dots,M\}$ (our tasks). A fixed adversary selects a stochastic opposing allocation strategy, i.e. a distribution $\pi_{\mathrm{adv}}$ over vectors $a=(a_{1},\dots,a_{M})$ which is fixed throughout training and evaluation. Let $s_{j}=\sum_{i=1}^{N}r_{ij}$ denote the team force committed to battlefield $j$. Our agents win against the adversary if the troops they allocate to a given field surpass the troops allocated by the adversary. The expected value secured on battlefield $j$ is therefore
$T_{j}(\mathbf{a}_{j})\;=\;v_{j}\,\mathbb{E}_{a\sim\pi_{\mathrm{adv}}}\!\bigl{[}\mathbf{1}\![\,s_{j}>a_{j}\,]\bigr{]}\;=\;v_{j}\,\Pr_{a\sim\pi_{\mathrm{adv}}}\!\bigl{[}s_{j}>a_{j}\bigr{]},$
where $1[x>y]$ denotes the indicator function. This is a thresholded-sum that remains symmetric and coordinate-wise non-decreasing in every agent’s contribution $r_{ij}$. Aggregating across battlefields with a value-weighted sum yields the team reward
$R(\mathbf{A})\;=\;\sum_{j=1}^{M}T_{j}(\mathbf{a}_{j})\;=\;\underbrace{\sum_{j=1}^{M}}_{U}\;T_{j}\!\Bigl{(}\underbrace{\sum_{i=1}^{N}}_{\operatorname*{\bigoplus}}r_{ij}\Bigr{)},$
so the game fits the double-aggregation structure $R(\mathbf{A})=\operatorname*{\bigoplus}_{j}\operatorname*{\bigoplus}_{i}r_{ij}$ assumed in our analysis.
Heterogeneity Gain: This is a continuous allocation game, and the inner aggregator $T_{j}$ is an indicator function over the sum of troop allocations to battlefield $j$. This function is Schur-concave (and Schur-convex at the same time!). Hence, by Thm. 3.2, heterogeneous colonel teams, where each colonel has a distinct troop allocation strategy, have no advantage over homogeneous teams, where all colonels employ the same allocation strategy: $\Delta R=0$. This makes sense, as it makes no difference whether two different colonels allocate $x/2$ troops to a battlefield, or one colonel allocates $x$ troops to the battlefield.
Our analysis also tells us what happens when we change $T_{j}$: this provides insights for generalizations of the Colonel Blotto game. For example, maybe the troops of different colonels don’t cooperate as well with each other, such that two colonels allocating $x/2$ troops to a battlefield results in a lower $T_{j}$-value than a single colonel allocating $x$ troops. In this case, $T_{j}$ becomes strictly Schur-convex, and Thm. 3.1 tells us that $\Delta R>0$ as long as the optimal allocation is non-trivial. Hence, heterogeneous teams are advantaged.
D.2 Level-Based Foraging
The well-known level-based foraging (LBF) benchmark, based on the knapsack problem [48], is a deceptively challenging, embodied MARL environment, where $N$ agents are placed on a grid with $M$ food items, and are tasked with collecting them. Each item $j$ has an integer level $L_{j}$ that must be met or exceeded by the combined skills of the agents standing on that cell before it can be collected [9]. Let agent $i$’s skill be $e_{i}$. At a given step the binary variable
$r_{ij}\;\in\;\{0,e_{i}\},\qquad\text{with}\;\sum_{j=1}^{M}r_{ij}\leq e_{i},$
denotes whether $i$ contributes its skill to item $j$. In our setting, we assume all agents are equally skilled, so $e_{i}=1\;\forall i$. Collecting these variables thus yields an allocation matrix $\mathbf{A}=[r_{ij}]\in\{0,e_{i}\}^{N\times M}$, which again matches our framework.
Inner aggregator.
A food item is harvested if the summed skill on its cell reaches the threshold, so
$T_{j}(\mathbf{a}_{j})\;=\;L_{j}\;\mathbf{1}\!\bigl{[}\textstyle\sum_{i=1}^{N}r_{ij}\;\geq\;L_{j}\bigr{]},\qquad\mathbf{a}_{j}=(r_{1j},\dots,r_{Nj})^{\top}.$
This threshold–sum is symmetric and monotone, depending only on the sum of its arguments and therefore simultaneously Schur-convex and Schur-concave.
Outer aggregator.
The stepwise team reward is the sum of harvested item values,
$R(\mathbf{A})\;=\;\sum_{j=1}^{M}T_{j}\!\Bigl{(}\sum_{i=1}^{N}r_{ij}\Bigr{)}\;=\;\underbrace{\sum_{j=1}^{M}}_{U}\;T_{j}\!\bigl{(}\underbrace{\sum_{i=1}^{N}}_{\operatorname*{\bigoplus}}r_{ij}\bigr{)}\;=\;\operatorname*{\bigoplus}_{j=1}^{M}\operatorname*{\bigoplus}_{i=1}^{N}r_{ij},$
so LBF also conforms to the double-aggregation form $R(\mathbf{A})=\operatorname*{\bigoplus}_{j}\operatorname*{\bigoplus}_{i}r_{ij}$.
MARL Environment Reward.
In the level-foraging environment, items that are picked up either disappear; replace themselves with different items; or replace themselves with the same item (possibly at a different cell). In all of these cases we can represent the cumulative reward as $\sum_{t=0}^{T}\gamma^{t}\,R_{t}\bigl{(}\mathbf{A}_{t}\bigr{)}$ for some sequence $(R_{t})_{t=1,\ldots T}$ of rewards adhering to the above reward structure.
Heterogeneity Gain: We analyze the heterogeneity gap of a specific stepwise reward $R$.
Because this is an embodied environment where each agent can either stand on an item ($r_{ij}=1$) or not ($r_{ij}=0$), effort allocations are discrete. Our continuous curvature test therefore does not apply directly, but the discrete analysis in Fig. 2 (left panel) does.
The table in Fig. 2 tells us something about the case where all items have level $L_{j}=1$. In this case, since we assumed $e_{i}=1$ for all agents, the inner aggregator reduces to
$T_{j}(\mathbf{a}_{j})=\mathbf{1}\!\Bigl{[}\textstyle\sum_{i=1}^{N}r_{ij}\geq 1\Bigr{]}=\max_{i}r_{ij},$
while the outer aggregator is an unnormalized sum, which becomes the mean when divided by $M$. Hence $R(\mathbf{A})=\sum_{j}\max_{i}r_{ij}$, which, up to the constant $1/M$, is exactly the case $U=\text{mean},\,T=\text{max}$ of Fig. 2. That table shows
$\frac{1}{M}\Delta R\;=\;\frac{\min\{M,N\}-1}{M},$
so the heterogeneity gap is strictly positive whenever the team could in principle cover more than one item ($\min\{M,N\}>1$). Intuitively, a homogeneous team can only collect one item per step (all agents flock to the same cell), whereas heterogeneous agents may spread out and capture up to $\min\{M,N\}$ items simultaneously.
This analysis can be extended to the case where all items have the same level $L>1$ and $L\mid N$ by grouping agents into $\tilde{N}:=N/L$ agent teams, each bundle contributing exactly $L$ units of skill. This yields
$\frac{1}{ML}\Delta R\;=\;\frac{\min\{M,\tilde{N}\}-1}{M}.$
(We omit the formal analysis, which is not difficult). Thus, if the team can form at least two such bundles ($\tilde{N}>1$), heterogeneity is again advantageous. If it cannot, then $\Delta R=0$, and there is no advantage to heterogeneity.
When the levels $\{L_{j}\}$ differ, an exact closed form is harder, but in general we expect $\Delta R>0$ whenever there is some combination of items that the heterogeneous team can collect, which in total is worth more than the largest single item that can be collected if all $N$ agents stand on its cell.
In LBF, therefore, our theory suggests that behavioral diversity is often advantageous. Note that (unlike the Colonel Blotto game) since LBF is an embodied, time-extended MARL environment, this analysis does not formally guarantee an advantage to RL-based heterogeneous agent teams: rather, it identifies that there are effort allocation strategies that will give these teams an advantage over homogeneous teams. The agents must still learn and be able to execute these strategies to gain this advantage (e.g., they must learn how to move to attain the desired allocations).
Appendix E Sum-Form Aggregators
Many useful reward functions are sum-form aggregators:
Definition E.1 (Sum-Form Aggregator).
A task-level aggregator $f:\mathbb{R}^{N}\to\mathbb{R}$ for task $j$ is a sum-form aggregator if it can be written as: $f(\mathbf{x}_{j})\;=\;\sum_{i=1}^{N}g(x_{j}),$ where $g_{j}:\mathbb{R}\to\mathbb{R}$ is differentiable. We say $f$ is (strictly) convex or concave if $g$ is (strictly) convex or concave, respectively.
Tab. 3 contains examples. When our aggregators have this form, Schur-convexity (concavity) is determined by whether $g$ is convex (concave)–a simple computational test. This is because of the following known connection between sum-form aggregators and Schur-convexity/concavity:
Lemma E.1 (Schur Properties of Sum-Form Aggregators [40]).
Given sum-form task-level aggregator $f(\mathbf{x})=\sum_{i=1}^{N}g(x_{i})$, the following holds: (i) if $g$ is (strictly) convex, then $f$ is (strictly) Schur-convex; and (ii) if $g$ is (strictly) concave, then $f$ is (strictly) Schur-concave.
This lemma simplifies checking the conditions of our heterogeneity gain results. For example, the following corollary can be used to establish $\Delta R>0$ for many of the aggregators in Tab. 3:
Corollary E.1 (Convex-Concave Positive Heterogeneity Gain).
Let $N,M\geq 2$. Let $g:[0,1]\to\mathbb{R}_{\geq 0}$ be a non-negative strictly convex function satisfying $g(0)=0$, and let $h:\mathbb{R}_{\geq 0}\to\mathbb{R}$ be a strictly concave, increasing function satisfying $h(0)=0$. If each task-level aggregator is a strictly convex sum-form aggregator $T_{j}(\mathbf{a}_{j})\;=\;\sum_{i=1}^{N}g\bigl{(}r_{ij}\bigr{)}$, and the outer aggregator is a strictly concave sum-form aggregator $U(\mathbf{y})=\sum_{j=1}^{M}h(y_{j})$, then $\Delta R>0$.
Proof of Corollary E.1.
We will apply Theorem 3.1 by verifying its conditions:
First, by Lemma E.1, since $g$ is strictly convex, the (identical) task-level aggregators $T_{j}(\mathbf{x})=\sum_{i=1}^{N}g(x_{i})$ are strictly Schur-convex, satisfying condition (i) of Theorem 3.1.
Second, the outer aggregator $U(y_{1},\ldots,y_{M})$ is strictly increasing at every coordinate by definition, satisfying condition (ii).
Hence, the conditions of Thm. 3.1 apply. To establish $\Delta R>0$, it remains to verify that the optimal allocation is non-trivial: it distributes effort across at least two tasks. In any admissible homogeneous solution, each of the $N$ agents chooses the same effort-distribution $(c_{1},\dots,c_{M})$ on tasks, with $\sum_{j}c_{j}=1$. Then task $j$’s reward is $T_{j}=N\,g(c_{j})$, so $R(\mathbf{A})\;=\;\sum_{j=1}^{M}h\!\bigl{(}N\,g(c_{j})\bigr{)}.$ The trivial, all-agent single-task allocation uses $(c_{j}=1,c_{k\neq j}=0)$. Its reward is therefore $R_{\text{corner}}=h\!\bigl{(}N\,g(1)\bigr{)}+\sum_{k\neq j}h\bigl{(}N\,g(0)\bigr{)}=h\!\bigl{(}N\,g(1)\bigr{)}$ since $g(0)=0$ and $h(0)=0$.
Strict concavity of $h$ implies that $h\!\bigl{(}N\,g(1)\bigr{)}<N\cdot h(g(1))$. Hence, agents can attain a better reward by allocating effort $1$ to $N$ different tasks rather than a single task. This shows that the best solution must use at least two nonzero $c_{j}$, completing the proof. ∎
Appendix F Formal Analysis
F.1 Proof of Thm. 3.1
Proof of Thm. 3.1.
Let $\mathbf{A}_{\mathrm{hom}}$ be an optimal homogeneous allocation (i.e., $R(A_{hom})=R_{hom}$), whose $i$th row is the vector
$\mathbf{c}\;=\;(c_{1},\dots,c_{M})\quad\text{with}\quad\sum_{j=1}^{M}c_{j}=1.$
Then each column $j$ of $\mathbf{A}_{\mathrm{hom}}$ is the uniform vector $\mathbf{u}_{j}\;=\;(c_{j},c_{j},\dots,c_{j})^{\top}\;\in\mathbb{R}^{N}.$ Hence the task-level reward is $T_{j}(\mathbf{u}_{j})$, and the overall reward is
$R\bigl{(}\mathbf{A}_{\mathrm{hom}}\bigr{)}=U\!\bigl{(}T_{1}(\mathbf{u}_{1}),\dots,T_{M}(\mathbf{u}_{M})\bigr{)}.$
Because $\sum_{j}c_{j}=1$, there is at least one task $j$ with $c_{j}>0$. We construct a heterogeneous allocation $A_{het}$ such that each column $x_{j}$ in $A_{het}$ has the same sum as the corresponding column in $A_{hom}$.
The total effort allocated to a task $j$ can be expressed as $\lfloor Nc_{j}\rfloor+f_{k}$, where $10\leq f_{j}<$ 1. First, we assign $\lfloor Nc_{j}\rfloor$ agents to allocate effort $1$ to task $j$, for every task $j$. These agents are all distinct. This leaves us with $\sum_{j}f_{j}~{}=~{}N~{}-~{}\sum_{j}\lfloor Nc_{j}\rfloor$ agents that have not allocated any effort yet. Let $i$ be the first of those agents. We have agent $i$ allocate $f_{1}$ effort to task $1$, $f_{2}$ effort to task $2$, and so on, until we arrive at a task $k$ such that $f_{1}+\ldots+f_{k}=1+s$, for some $s>0$. We have $i$ allocate $f_{k}-s$ to this task $k$. Then, we move to agent $i+1$, and allocate the remaining fractional efforts in the same manner (and in particular, allocating $s$ effort to task $k$), until agent $i+1$ overflows. Then we move to agent $i+2$, and so on. This ensures that we have allocated $N$ effort in total across the agents, and that every agent’s effort allocation sums exactly to $1$, so is feasible.
Let $x_{j}$ be the $j$th column of $A_{het}$. We note the following fact: any non-uniform vector whose sum is $Nc_{j}$ majorizes the uniform vector $u_{j}$. Hence, $T_{j}(x_{j})\geq T_{j}(u_{j})$, with equality only if $x_{j}=u_{j}$. This means that if $A_{het}\neq A_{hom}$, then
$R\bigl{(}\mathbf{A}_{\mathrm{het}}\bigr{)}\;=\;U\!\bigl{(}T_{1}(\mathbf{x}_{1}),\dots,T_{M}(\mathbf{x}_{M})\bigr{)}\;>\;U\!\bigl{(}T_{1}(\mathbf{u}_{1}),\dots,T_{M}(\mathbf{u}_{M})\bigr{)}\;=\;R\bigl{(}\mathbf{A}_{\mathrm{hom}}\bigr{)}.$
We note that $A_{hom}=A_{het}$ only if $A_{hom}$ is a trivial allocation, as $A_{het}$ contains at least one agent allocating effort $1$ to some task, and $A_{hom}$’s agents only allocate fractional efforts, if it is non-trivial. Otherwise, since $R(A_{hom})=R_{hom}$, the above inequality implies $\Delta R=R_{\mathrm{het}}-R_{\mathrm{hom}}>0$. This completes the proof. ∎
F.2 Proof of Thm. 3.2
Proof of Thm. 3.2.
Let $\mathbf{A}$ be an arbitrary feasible allocation, and let $\mathbf{A}_{\mathrm{hom}}$ be a homogeneous allocation with the same column sums. Concretely, for each column $j$, define
$s_{j}\;=\;\sum_{i=1}^{N}r_{ij}\quad\text{and}\quad\mathbf{u}_{j}\;=\;\bigl{(}\tfrac{s_{j}}{N},\tfrac{s_{j}}{N},\dots,\tfrac{s_{j}}{N}\bigr{)}^{\top},$
so $\mathbf{u}_{j}$ is the uniform distribution of total mass $s_{j}$ across $N$ agents. Then construct
$\mathbf{A}_{\mathrm{hom}}\;=\;\begin{pmatrix}\tfrac{s_{1}}{N}&\cdots&\tfrac{s_{M}}{N}\\ \vdots&\ddots&\vdots\\ \tfrac{s_{1}}{N}&\cdots&\tfrac{s_{M}}{N}\end{pmatrix},$
which is clearly homogeneous (each row is the same), and respects each column sum $s_{j}$. Since $\sum_{j}s_{j}=N$, each row sums to $1$, hence the allocation is feasible. By Schur-concavity of $T_{j}$, for each column $j$ we have
$\mathbf{a}_{j}\;\succ\;\mathbf{u}_{j}\quad\Longrightarrow\quad T_{j}(\mathbf{a}_{j})\;\leq\;T_{j}(\mathbf{u}_{j}),$
unless $\mathbf{a}_{j}$ is $\mathbf{u}_{j}$. In other words, any deviation from the uniform vector with the same sum $\sum_{i=1}^{N}a_{ji}=s_{j}$ will not increase $T_{j}(\mathbf{a}_{j})$ under Schur-concavity. Hence for each column $j$ of $\mathbf{A}$, $T_{j}(\mathbf{a}_{j})\;\leq\;T_{j}(\mathbf{u}_{j}),$. Since $U$ is non-decreasing in each coordinate,
$R\bigl{(}\mathbf{A}\bigr{)}\;=\;U\!\bigl{(}T_{1}(\mathbf{a}_{1}),\dots,T_{M}(\mathbf{a}_{M})\bigr{)}\;\leq\;U\!\bigl{(}T_{1}(\mathbf{u}_{1}),\dots,T_{M}(\mathbf{u}_{M})\bigr{)}\;=\;R\bigl{(}\mathbf{A}_{\mathrm{hom}}\bigr{)}.$
[…]
F.3 Proof of Thm. 3.3
Proof of Thm. 3.3.
By hypothesis, the components of the task score vector
$\textbf{T}(\textbf{A})=\Bigl{(}T_{1}(\mathbf{a}_{1}),\,T_{2}(\mathbf{a}_{2}),\,\dots,\,T_{M}(\mathbf{a}_{M})\Bigr{)}$
[…]
Let each agent $i$ invest all of its effort into task 1. This is the trivial allocation. Then the first column of $\mathbf{A}_{\mathrm{corner}}$ is $(1,1,\dots,1)^{\top}$, and all other columns $\mathbf{a}_{j}$ are zero. Since task scores sum to $C$, we get $T_{1}\bigl{(}\mathbf{a}_{1}\bigr{)}=C,\quad T_{j}\bigl{(}\mathbf{a}_{j}\bigr{)}=0\;\;\text{for }j\neq 1.$ By assumption (2), we infer that the vector of task-level scores is indeed $(C,0,\dots,0)$.
Notice that each row of $A_{corner}$ is the same $(1,0,\dots,0)$, making $\mathbf{A}_{\mathrm{corner}}$ a homogeneous allocation. Hence, we attained the maximum possible reward $R(\textbf{A})$ through a homogeneous allocation, implying $\Delta R=0$. ∎
F.4 Proof of Thm. 3.4
Before proving the statement, let’s write the expressions for homogeneous and heterogeneous optima. For each task $j$, we defined
$T_{j}(\mathbf{A})=\sum_{i=1}^{N}\frac{\exp\bigl{(}t\cdot r_{ij}\bigr{)}}{\sum_{\ell=1}^{N}\exp\bigl{(}t\cdot r_{\ell j}\bigr{)}}\;r_{ij},$
while defining the outer aggregator to be
$U\bigl{(}T_{1}(\mathbf{a}_{1}),\ldots T_{M}(\mathbf{a}_{m}))=\sum_{j=1}^{M}\frac{\exp\bigl{(}\tau\cdot T_{j}(\mathbf{A})\bigr{)}}{\sum_{\ell=1}^{M}\exp\bigl{(}\tau\cdot T_{\ell}(\mathbf{A})\bigr{)}}\;T_{j}(\mathbf{A}),$
where $t,\tau\in\mathbb{R}$ are temperature parameters. In the homogeneous setting, where all agents share the same allocation $\mathbf{c}=(c_{1},\dots,c_{M})$, we therefore have $T_{j}(\mathbf{A})=\sum_{i=1}^{N}\frac{\exp\bigl{(}t\,c_{j}\bigr{)}}{\sum_{\ell=1}^{N}\exp\bigl{(}t\,c_{j}\bigr{)}}\;c_{j}=c_{j}.$ Thus,
$R_{\mathrm{hom}}=\max_{\mathbf{c}\,\in\,\Delta^{M-1}}\quad\sum_{j=1}^{M}\frac{\exp\bigl{(}\tau\,c_{j}\bigr{)}}{\sum_{\ell=1}^{M}\exp\bigl{(}\tau\,c_{\ell}\bigr{)}}\;c_{j}$
where $\Delta^{M-1}$ is the simplex of all admissible allocations.
In the general heterogeneous setting, each row $(r_{i1},\dots,r_{iM})$ can be different. Then
$T_{j}(\mathbf{A})\;=\;\sum_{i=1}^{N}\frac{\exp\bigl{(}t\,r_{ij}\bigr{)}}{\sum_{\ell=1}^{N}\exp\bigl{(}t\,r_{\ell j}\bigr{)}}\;r_{ij},$
and we choose $\mathbf{A}\in(\Delta^{M-1})^{N}$ to maximize
$R_{\mathrm{het}}\;=\;\max_{\mathbf{A}}\sum_{j=1}^{M}\frac{\exp\bigl{(}\tau\,T_{j}(\mathbf{A})\bigr{)}}{\sum_{k=1}^{M}\exp\bigl{(}\tau\,T_{k}(\mathbf{A})\bigr{)}}\;T_{j}(\mathbf{A}).$
Keeping these expressions in mind, we proceed with the proof of Thm. 3.4.
Reminder: assuming $N=M\geq 2$, we want to prove $\Delta R(t,\tau;N)=0$ when $t\leq 0$, and
[…]
otherwise, where $\sigma(t,N)\;:=\;\frac{e^{t}}{e^{t}+N-1}.$
Proof of Thm. 3.4.
When $t\leq 0$, $T_{j}$ is Schur–concave, so $\Delta R=0$ by Thm. 3.2. We assume $t>0$ for the rest of the proof.
Homogeneous optimum. If every row of $\mathbf{A}$ equals the same allocation $\mathbf{c}\in\Delta^{N-1}$, then $T_{j}(\mathbf{A})=c_{j}$. $U$ is Schur–concave for $\tau\leq 0$, and Schur–convex for $\tau\geq 0$, hence it is maximized by the uniform distribution in the former case, and by a $1$-hot vector in the latter case, yielding:
$R_{\mathrm{hom}}\;=\;\max_{\mathbf{c}\in\Delta^{N-1}}U(\mathbf{c})\;=\;\begin{cases}\dfrac{1}{N},&\tau\leq 0,\\[6.0pt] \sigma(\tau,N),&\tau>0.\end{cases}$
Lower bound on $R_{het}$. The trivial allocation, where every agent works on the same task, produces $R_{\mathrm{trivial}}=\sigma(\tau,N)$. The spread allocation, where agent $i$ works exclusively on task $i$, makes each column “one-hot”; this gives $T_{j}=\sigma(t,N)$ for all $j$, and plugging this into $U$, we get $R_{\mathrm{spread}}=\sigma(t,N)$. Consequently
$R_{\mathrm{het}}\;\geq\;\max\{\sigma(t,N),\,\sigma(\tau,N)\}.$
Combining (H) and (L) gives the desired lower bound. ∎
Appendix G Deriving the $\{\min,\textnormal{mean},\max\}$ heterogeneity gains in the Fig. 2 table
We derive these heterogeneity gain case-by-case. Tab. 1 summarizes the derivation for continuous allocations ($r_{ij}\in[0,1]$), and Tab. 2 does the same for discrete effort allocations ($r_{ij}\in\{0,1\}$).
Table 1: All nine extreme cases of inner/outer aggregators belonging to the set $\{\min,\text{mean},\max\}$. In each cell, we show the best possible outcome for Heterogeneous vs. Homogeneous allocations and the resulting $\Delta R$.
$T=\min$ $T=\text{mean}$ $T=\max$
$U=\min$ Inner: $T_{j}=\min_{i}r_{ij}.$ Best $R_{het}$, $R_{hom}$: All must have $r_{ij}\geq x$ to push $\min_{i}r_{ij}\!=\!x$, so $x\leq 1/M.$ $\implies T_{j}=1/M.$ Outer: $\min_{j}T_{j}=1/M\implies R=1/M.$ Gap: 0. Inner: $T_{j}=\tfrac{1}{N}\sum_{i}r_{ij}$ (avg over $i$). Maximize $\min_{j}T_{j}$: Both $R_{het}$, $R_{hom}$ must make $T_{j}$ all equal (for best min), so $T_{j}=1/M.$ Outer: $\min_{j}T_{j}=1/M\implies R=1/M.$ Gap: 0. Inner: $T_{j}=\max_{i}r_{ij}.$ Outer: picks $\min_{j}T_{j}$. $R_{het}$: $\min_{j}T_{j}=1\implies R=1.$ $R_{hom}$: $\min_{j}T_{j}=1/M\implies R=1/M.$ Gap: $1-\tfrac{1}{M}=\tfrac{M-1}{M}.$
$U=$ mean Inner: $T_{j}=\min_{i}r_{ij}=1/M.$ Outer: simple avg $\tfrac{1}{M}\sum_{j}T_{j}.$ Since $\sum_{j}T_{j}=M\cdot(1/M)=1\implies R=1/M.$ Both $R_{het}$, $R_{hom}$ same $\implies\Delta R=0.$ Gap: 0. Inner: $T_{j}=\tfrac{1}{N}\sum_{i}r_{ij}.$ Then $\sum_{j}T_{j}=1.$ Outer: avg $\!=\!\tfrac{1}{M}\sum_{j}T_{j}.$ Hence $R=\tfrac{1}{M}\cdot 1=\tfrac{1}{M}.$ Same for $R_{het}$, $R_{hom}$. Gap: 0. Inner: $T_{j}=\max_{i}r_{ij}.$ Outer: avg $\!=\!\tfrac{1}{M}\sum_{j}T_{j}.$ $R_{het}$: sum $=M\implies R=1.$ $R_{hom}$: sum $=1\implies R=1/M.$ Gap: $1-\tfrac{1}{M}=\tfrac{M-1}{M}.$
$U=\max$ Inner: $T_{j}=\min_{i}r_{ij}$ can be made $1$ for one task. Outer: picks $\max_{j}T_{j}=1\implies R=1$ Same for $R_{het}$, $R_{hom}$. Gap: 0. Inner: $T_{j}=$ avg over $i.$ Outer: picks $\max_{j}T_{j}.$ Both $R_{het}$, $R_{hom}$ can put all effort into one task to get $T_{j}=1,$ so $R=1.$ Gap: 0. Inner: $T_{j}=\max_{i}r_{ij}.$ Outer: picks $\max_{j}T_{j}.$ Both $R_{het}$, $R_{hom}$ can achieve $\max_{j}=1\implies R=1.$ Gap: 0.
Table 2: A “9 extreme cases” table for discrete, one-task-per-agent allocations.
$\min$ mean $\max$
$\min$ Inner: $T_{j}\!\to\!\begin{cases}1,&\text{if all agents pick $j$},\\ 0,&\text{otherwise}.\end{cases}$ Outer: $\min_{j}T_{j}$. To get $R>0$, must have $T_{j}>0$ for every $j$ (i.e. all agents pick all tasks, impossible). Hence $R_{\mathrm{het}}=R_{\mathrm{hom}}=0$ typically, $\Delta R=0.$ Inner: $T_{j}=\frac{|\mathcal{I}_{j}|}{N}$ Outer: $\min_{j}T_{j}$. $R_{het}=\lfloor N/M\rfloor/N$. $R_{hom}=0$. $\Delta R=\lfloor N/M\rfloor/N$. Inner: $T_{j}\!\to\!\begin{cases}1,&\text{if at least 1 agent picks $j$},\\ 0,&\text{if no agent picks $j$}.\end{cases}$ Outer: $\min_{j}T_{j}.$ Heterogeneous can choose $s$ distinct tasks. If want $\min_{j}=1$, must pick all $M$ tasks. That requires $N\geq M$. Then $R=1.$ Homogeneous covers only 1 task $\implies\min_{j}=0$ for $M>1\implies R=0.$ $\Delta R=1$ if $N\geq M$, else $0$
mean Inner: $T_{j}=1$ only if all pick $j$, else 0. Summation $\sum_{j}T_{j}$ is number of tasks chosen by all agents. Usually 0 or 1. Outer: Average across $j$. $R=\frac{1}{M}\sum_{j}T_{j}.$ $\implies R=1/M,$ $\Delta R=0.$ Inner: $T_{j}=\tfrac{|\mathcal{I}_{j}|}{N}$. Outer: Average across tasks: $R=\frac{1}{M}\sum_{j=1}^{M}\frac{|\mathcal{I}_{j}|}{N}=\frac{1}{M}.$ No matter how agents are distributed, $\sum_{j=1}^{M}|\mathcal{I}_{j}|=N$. Hence $R_{\mathrm{het}}=R_{\mathrm{hom}}=\tfrac{1}{M},$ $\Delta R=0.$ Inner: $T_{j}=1$ if chosen by at least 1 agent, else 0. Outer: Average across $j$: $\tfrac{1}{M}\sum_{j}T_{j}.$ This is $\tfrac{1}{M}\cdot\bigl{(}\#\text{ of tasks chosen}\bigr{)}.$ Heterogeneous can pick up to $\min(M,N)$ tasks, so $R=\frac{\min(M,N)}{M}.$ Homogeneous covers exactly 1 task $\implies R=1/M.$ $\Delta R=\frac{\min(M,N)-1}{M}.$
$\max$ Inner: $T_{j}=1$ only if all pick $j$, else 0. Outer: $\max_{j}T_{j}.$ $\Delta R=0.$ Inner: $T_{j}=|\mathcal{I}_{j}|/N.$ Outer ($\tau\to+\infty$): $\max_{j}T_{j}$. We can place all agents on one task, get $T_{j}=1.$ Then $R=1.$ Same for homogeneous or heterogeneous. $\Delta R=0.$ Inner: $T_{j}=1$ if at least 1 picks $j$, else 0. Outer: $\max_{j}T_{j}=1$ if any agent picks $j$. Even a single task yields $R=1.$ So $R_{\mathrm{hom}}=R_{\mathrm{het}}=1,$ $\Delta R=0.$
Appendix H Parametrized Families of Aggregators
The Table in this section illustrates several families of generalized aggregators that the analysis in this paper applies to. The scalar $t$ parametrizes each family of aggregators, continuously shifting the aggregators from Schur-concave to Schur-convex.
Table 3: Illustrative families of parametric (and one nonparametric) aggregators $f_{t}(\mathbf{x})$. Changing the real parameter $t$ can switch between Schur-convex and Schur-concave behaviors (on nonnegative inputs), or control how strongly the aggregator favors “peaked” vs. “uniform” distributions. As $t\to\pm\infty$ or $t\to 0$, many reduce to well-known extremes such as $\max$, $\min$, or the arithmetic mean.
Name Definition Schur Property & Limits
Power-Sum $f_{t}(\mathbf{x})~{}=~{}\sum_{i=1}^{N}\bigl{(}x_{i}\bigr{)}^{t},\quad x_{i}\geq 0,\quad t>0$ Strictly Schur-convex for $t>1$. Strictly Schur-concave for $10<t<$ 1. At $t=1$, it is linear (both Schur-convex and Schur-concave). Undefined at $t\leq 0$ if any $x_{i}=0$, though one can extend with limits.
Power-Mean $M_{t}(\mathbf{x})~{}=~{}\biggl{(}\frac{1}{N}\sum_{i=1}^{N}(x_{i})^{t}\biggr{)}^{\!\!1/t},\quad x_{i}\geq 0,\;t\neq 0$ Strictly Schur-convex for $t>1$. Strictly Schur-concave for $10<t<$ 1. Reduces to arithmetic mean at $t=1$. As $t\to\infty$, converges to $\max_{i}x_{i}$; as $t\to-\infty$, converges to $\min_{i}x_{i}$.
Log-Sum-Exp (LSE) $\mathrm{LSE}_{t}(\mathbf{x})~{}=~{}\frac{1}{t}\ln\!\Bigl{(}\sum_{i=1}^{N}e^{\,t\,x_{i}}\Bigr{)},\quad t\neq 0$ Strictly Schur-convex for $t>0$. Strictly Schur-concave for $t<0$. As $t\to\infty$, approaches $\max_{i}x_{i}$; as $t\to-\infty$, approaches $\min_{i}x_{i}$.
Softmax Aggregator $\mathrm{Softmax}_{t}(\mathbf{x})~{}=~{}\sum_{i=1}^{N}\frac{e^{\,t\,x_{i}}}{\sum_{j=1}^{N}e^{\,t\,x_{j}}}\,x_{i},\quad t\in\mathbb{R}$ Strictly Schur-convex for $t>0$. Strictly Schur-concave for $t<0$. As $t\to\infty$, converges to $\max_{i}x_{i}$; as $t\to-\infty$, converges to $\min_{i}x_{i}$. At $t=0$, each weight is $\tfrac{1}{N}$, so $\mathrm{Softmax}_{0}(\mathbf{x})=\tfrac{1}{N}\sum_{i}x_{i}$.
Appendix I Additional Results in the Multi-Agent Multi-Task Matrix Game
We report further details and results on the heterogeneity gains obtained in the multi-agent multi-task matrix game.
I.1 Game formulation
In Tab. 4 we provide an example of the pay-off matrix in this game for $N=M=3$.
Table 4: Example of a Multi-Agent Multi-Task matrix game for $N=M=3$. Agents choose their actions $A=(r_{ij})$ and receive the global reward $R(\mathbf{A})=\operatorname*{\bigoplus}_{j=1}^{M}\operatorname*{\bigoplus}_{i=1}^{N}r_{ij}$.
Tasks
1 2 3
Agents 1 $r_{11}$ $r_{12}$ $r_{13}$
2 $r_{21}$ $r_{22}$ $r_{23}$
3 $r_{31}$ $r_{32}$ $r_{33}$
I.2 $N=M=2$
We train with $N=2,M=2$ for 100 training iterations (each consisting of 60,000 frames). We report the results for the continuous case in Tab. 5 and for the discrete case in Tab. 6. The evolution of the heterogeneity gains over training is shown in Fig. 6.
Table 5: Heterogeneity gain $\Delta R\in\mathbb{R}_{0\leqslant x\leqslant 1}$ of the continuous matrix game with $N=M=2$. The results match the theoretical analysis in the Table of Fig. 2. We report mean and standard deviation after 6 million training frames over 9 different random seeds.
$T$
Min Mean Max
$U$ Min $-0.002\pm 0.002$ $0.000\pm 0.003$ $0.504\pm 0.007$
Mean $-0.002\pm 0.002$ $000.000\pm 0$ 0.000 $0.496\pm 0.001$
Max $-0.003\pm 0.002$ $-0.001\pm 0.001$ $0.003\pm 0.001$
Table 6: Heterogeneity gain $\Delta R\in\mathbb{R}_{0\leqslant x\leqslant 1}$ of the discrete matrix game with $N=M=2$. The results match the theoretical analysis in Fig. 2. We report mean and standard deviation after 6 million training frames over 9 different random seeds.
$T$
Min Mean Max
$U$ Min $0.0\pm 0$ 0.0 $0.5\pm 0.0$ $1\pm 0.0$
Mean $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.5\pm 0.0$
Max $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
Figure 6: Heterogeneity gain for the discrete and continuous matrix games with $N=M=2$ over training iterations. We report mean and standard deviation after 6 million training frames over 9 different random seeds. The final results match the theoretical predictions in Fig. 2.
I.3 $N=M=4$
In the case $N=M=4$, the evolution of the heterogeneity gains during training is shown in Fig. 3. We further report the final obtained gains for the continuous case in Tab. 7 and for the discrete case in Tab. 8.
Table 7: Heterogeneity gain $\Delta R\in\mathbb{R}_{0\leqslant x\leqslant 1}$ of the continuous matrix game with $N=M=4$. The results match the theoretical analysis in Fig. 2. We report mean and standard deviation after 12 million training frames over 9 different random seeds.
$T$
Min Mean Max
$U$ Min $-0.003\pm 0.002$ $0.000\pm 0.001$ $0.690\pm 0.026$
Mean $-0.002\pm 0.000$ $000.000\pm 0$ 0.000 $0.722\pm 0.002$
Max $-0.037\pm 0.023$ $-0.009\pm 0.005$ $0.029\pm 0.006$
Table 8: Heterogeneity gain $\Delta R\in\mathbb{R}_{0\leqslant x\leqslant 1}$ of the discrete matrix game with $N=M=4$. The results match the theoretical analysis in the Table of Fig. 2. We report mean and standard deviation after 12 million training frames over 9 different random seeds.
$T$
Min Mean Max
$U$ Min $0.0\pm 0$ 0.0 $0.25\pm 0.0$ $1.0\pm 0.0$
Mean $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.75\pm 0.0$
Max $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
Appendix J Multi-Goal-Capture
In Fig. 7 we juxtapose two representative $N=M=2$ roll-outs of the Multi-Goal-Capture environment for homogeneous teams (top row) and heterogeneous teams (bottom row) when $U~{}=~{}\min,\,T=\max$. Consistent with the discussion in Sec. 5, homogeneous agents steer to the geometric midpoint between the two goals, producing almost overlapping paths–this is suboptimal, as they cannot cover both goals. On the other hand, heterogeneous agents exaggerate their differences–taking sharply diverging trajectories and ensuring one goal each.
(a) Homogeneous run 1
(b) Homogeneous run 2
(c) Heterogeneous run 1
(d) Heterogeneous run 2
Figure 7: Behaviour under the concave–convex aggregator $U=\min,\,T=\max$. Each dot is an agent position; line segments indicate instantaneous velocity; green squares mark goal locations. Homogeneous policies collapse to a single “mid-point” route, while heterogeneous policies split and follow distinct paths to cover both goals. Note how the heterogeneous agents exaggerate the difference in their trajectories, rather than head directly to the goal: this is an outcome of the reward structure, which encourages maximal diversity.
Appendix K Parametrized Dec-POMDP
A Parametrized Decentralized Partially Observable Markov Decision Process (PDec-POMDP) is defined as a tuple
$\left\langle\mathcal{N},\mathcal{S},\left\{\mathcal{O}_{i}\right\}_{i\in\mathcal{N}},\left\{\sigma_{i}^{\theta}\right\}_{i\in\mathcal{N}},\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}},\mathcal{R}^{\theta},\mathcal{T}^{\theta},\gamma,s_{0}^{\theta}\right\rangle_{\theta},$
where $\mathcal{N}=\{1,\ldots,n\}$ denotes the set of agents, $\mathcal{S}$ is the state space, and, $\left\{\mathcal{O}_{i}\right\}_{i\in\mathcal{N}}$ and $\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}$ are the observation and action spaces, with $\mathcal{O}_{i}\subseteq\mathcal{S},\;\forall i\in\mathcal{N}$. Further, $\left\{\sigma_{i}^{\theta}\right\}_{i\in\mathcal{N}}$ and $\mathcal{R}^{\theta}$ are the agent observation and reward functions, such that $\sigma_{i}^{\theta}:\mathcal{S}\mapsto\mathcal{O}_{i}$, and, $\mathcal{R}^{\theta}:\mathcal{S}\times\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}\mapsto\mathbb{R}$. $\mathcal{T}^{\theta}$ is the stochastic state transition model, defined as $\mathcal{T}^{\theta}:\mathcal{S}\times\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}\mapsto\Delta\mathcal{S}$, which outputs the probability $\mathcal{T}^{\theta}(s^{t},\left\{a^{t}_{i}\right\}_{i\in\mathcal{N}},s^{t+1})$ of transitioning to state $s^{t+1}\in\mathcal{S}$ given the current state $s^{t}\in\mathcal{S}$ and actions $\left\{a^{t}_{i}\right\}_{i\in\mathcal{N}}$, with $a^{t}_{i}\in\mathcal{A}_{i}$. $\gamma$ is the discount factor. Finally, $s_{0}^{\theta}\in\mathcal{S}$ is a the initial environment state. A PDec-POMDP represents a set of traditional Dec-POMDPs [41], where the observation function, the transition function, the reward function, and the initial state are conditioned on parameters $\theta$. This formalism is similar to the concepts of Underspecified POMDP [36] and contextual MDP [49].
Agents are equipped with (possibly stochastic) policies $\pi_{i}(a_{i}|o_{i})$, which compute an action given a local observation. Their objective is to maximize the discounted return:
$G^{\theta}(\mathbf{\pi})=\mathbb{E}_{\mathbf{\pi}}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}^{\theta}\left(s^{t},\mathbf{a}^{t}\right)\middle|s^{t+1}\sim\mathcal{T}^{\theta}(s^{t},\mathbf{a}^{t}),a_{i}^{t}\sim\pi_{i}(o_{i}^{t}),o_{i}^{t}=\sigma^{\theta}_{i}(s_{t})\right],$
where $\mathbf{\pi},\mathbf{a}$ are the vectors of all agents’ policies and actions. $G^{\theta}(\mathbf{\pi})$ represents the expected sum of discounted rewards starting in state $s_{0}^{\theta}$ and following policy $\mathbf{\pi}$ in a PDec-POMDP parametrized by $\theta$.
Appendix L Limitations and Open Questions
We list a number of limitations, open questions, and possible extensions.
L.1 Theoretical scope
  • Inseparable tasks. This work considers task allocation problems where tasks are separable (i.e., the subtasks are independent), and that team reward can be factorized as an inner symmetric aggregator followed by an outer symmetric aggregator. Extending the theory to inseparable rewards, where performance on one task causally affects performance on another task (e.g., coverage with overlap penalty, or pairwise task constraints), is an open problem.
  • Beyond task-allocation RL domains. While the benchmark domains we study (matrix game, multi-goal capture) and the additional settings covered in App. D (Colonel Blotto, level-based foraging) all fit naturally into the task-allocation template, many notable RL domains—team autonomous driving, multi-robot manipulation—might not be representable within a task allocation framework. Our heterogeneity analysis does not directly apply to these settings, and extending our results to them is important for getting a complete picture of the benefits of heterogeneity.
L.2 Algorithmic assumptions
  • Differentiable simulation. HED requires $\nabla_{\theta}G^{\theta}(\bm{\pi})$, hence a simulator that is end-to-end differentiable and tractable to back-propagate through. Many realistic environments still rely on non-smooth physics or black-box generators, requiring us to modify HED for these settings. We note that there are good methods for learning environment parameters in non-differentiable settings, including the PAIRED algorithm [36] and the MARL-based environment design algorithm of [35], which can potentially be used to extend HED—but we leave such extensions to future work.
L.3 Open questions
  1. i.
    What is the connection between the transition function and heterogeneity? Our analysis is reward-centric: the curvature criterion reasons only about the team reward. In a Dec-POMDP, however, heterogeneity can be beneficial purely because agents are constrained by state transitions. When do state transition dynamics benefit heterogeneity?
  2. ii.
    Learning dynamics vs. reward structure. The theory predicts whether a given reward structure enables an advantage to heterogeneous teams, not whether a particular learning algorithm will learn in response to it. Our experiments suggest, empirically, that heterogeneous agents will, in practice, learn to exploit heterogeneous reward structures; but can a formal link be established between our reward structure insights and what reward the learning dynamics converge to in practice?
Tackling these challenges would sharpen our understanding of when and how diversity should be engineered in cooperative multi-agent learning.